OpenStudy (anonymous):

Solve for x: 2/5x + 2 = 10 x = 20 x = 24 x = 4 x = 8

OpenStudy (anonymous):

(2/5)x = 10-2 (2/5)x = 8 x = 8(5/2) x = 20 question ? :D